OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (eyust707):

Well if two lines are parallel they both have the same slope. What is the slope of y=-x+6 ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

*-1

OpenStudy (eyust707):

yes!

OpenStudy (eyust707):

So we know the slope of the line we are about to create is -1

OpenStudy (eyust707):

The standard form of a line looks like this: \[Ax+ By = C\]

OpenStudy (anonymous):

ax+by=-1?

OpenStudy (eyust707):

but that doesnt really help us cause I don't see slope anywhere.... We have slope and we have a point, so lets use point slope form, then switch back to the form they want. Point slope form is \[(y - y_1) = m (x - x_1)\]

OpenStudy (eyust707):

y1 is given x1 is given and now we know m

OpenStudy (eyust707):

Thats all we need.. just plug those in and move stuff around until it is in standard form